OpenStudy (anonymous):

What is the common ratio of the geometric sequence whose second and fourth terms are 6 and 54, respectively? A:1 B:2 C;3 D;4

OpenStudy (noelgreco):

You must be studying this stuff, what have you tried?

jimthompson5910 (jim_thompson5910):

a2 is the second term r is the common ratio a3 is the third term a3 = a2*r a3 = 6*r agreed so far?